# Just Shift It: Test-Time Prototype Shifting for Zero-Shot Generalization with Vision-Language Models

### v1

@@ 4.3.3 Cross-Dataset Generalization
Ours (Shift + templates + descriptors) 71.54${}_{(\uparrow{4.26})}$ 50.47${}_{(\uparrow{6.03})}$ 87.35${}_{(\downarrow{0.63})}$ 69.06${}_{(\uparrow{3.82})}$ 71.00${}_{(\uparrow{5.92})}$ 95.09${}_{(\uparrow{2.11})}$ 85.23${}_{(\uparrow{1.43})}$ 68.98${}_{(\uparrow{6.43})}$ 26.34${}_{(\uparrow{2.64})}$ 44.48${}_{(\uparrow{3.06})}$ 66.96${}_{(\uparrow{3.51})}$
@@
4.3.4 Comparison to Training-Free Methods
SuS-X-SD-C [47]
61.65
61.69
35.88
60.49
SuS-X-LC-P [47]
61.80
61.62
36.25
60.64
CALIP [11]
60.57
59.34
Ours (Shift + SuS-X descriptors) 63.52 63.66 37.66 61.47
@@ 4.5.1 Effect of Shift on Different Prototypes
Table 5: Acc@1 for zero-shot image classification comparing pure zero-shot baselines vs. with learned feature-space shift from prototypes derived from various prompts using CLIP-ViT/B-16 backbone.
Prompt Type Setting ImageNet ImageNet OOD Average Cross-Dataset Average
Vanilla
Zero-Shot
66.74
57.20
63.45
 + shift
68.77
61.59
64.41
$\mathbf{\Delta}$
 + 2.03  + 4.39  + 0.96
CoOp [56]
Zero-Shot
71.51
59.28
N/A
 + shift
73.73
63.46
N/A
$\mathbf{\Delta}$
 + 2.22  + 4.18 N/A
CLIP templates
Zero-Shot
68.35
59.43
64.69
 + shift
70.38
64.04
65.57
$\mathbf{\Delta}$
 + 2.03  + 4.61  + 0.88
Descriptors
Zero-Shot
68.52
58.29
66.02
 + shift
70.40
62.48
66.80
$\mathbf{\Delta}$
 + 1.88  + 4.19  + 0.78
CLIP templates + Descriptors
Zero-Shot
69.54
59.88
65.94
 + shift
71.45
64.15
66.96
$\mathbf{\Delta}$
 + 1.91  + 4.27  + 1.02
@@ 4.5.2 Feature-Space Transformation Variants
Method ImageNet ImageNet OOD Average Cross-Dataset Average
Scale
70.40
61.12
65.96
@@ 5 Conclusion
In this work, we have presented the Test-time Prototype Shifting (TPS) framework, a novel approach to enhancing the zero-shot generalization abilities of VLMs. TPS addresses the limitations of existing test-time training methods, such as Text-Prompt Tuning, by directly modulating class prototypes in the embedding space. This strategy not only reduces the computational and memory demands significantly but also allows for greater flexibility and precision in adapting to diverse domain shifts. By leveraging pre-computed and cached prototypes, and introducing class-specific shifts, TPS effectively bridges domain gaps. The extensive evaluations conducted across multiple datasets demonstrate the superior performance of our method, outperforming existing approaches in terms of accuracy while being more computationally efficient. This work not only sets a new standard for zero-shot generalization in VLMs but also presents a scalable, efficient solution for real-world applications.
6 Research Impact and Limitations
Appendix 0.A Implementation Details of TPS
@@
Appendix 0.B Main Results With More Random Seeds
In Sec 0.B.1 and  0.B.2, we run Test-Time Prototype Shifting (TPS) over 3 random seeds on both the natural distribution shifts (Table 1) and cross-dataset generalization (Table 3), respectively. The randomness comes from the image augmentation in creating a diverse minibatch for the entropy minimization objective.
0.B.1 Natural Distribution Shifts
@@
Method ImageNet ImageNet-A ImageNet-V2 ImageNet-R ImageNet-Sketch Average
OOD Average
Test-Time Adaptation Baselines
TPT [43]
68.96 (±.$(\pm{.03})$
54.47 (±.$(\pm{.26})$
63.46 (±.$(\pm{.07})$
77.10 (±.$(\pm{.04})$
47.93 (±.$(\pm{.03})$
62.38 (±.$(\pm{.05})$
60.74 (±.$(\pm{.06})$
TPT + (templates + descriptors)*
69.51 (±.$(\pm{.05})$
54.94 (±.$(\pm{.17})$
63.86 (±.$(\pm{.11})$
77.57 (±.$(\pm{.11})$
48.38 (±.$(\pm{.04})$
62.85 (±.$(\pm{.03})$
61.19 (±.$(\pm{.04})$
Ours
71.43 (±.$(\pm{.06})$
60.78 (±.$(\pm{.21})$
65.00 (±.$(\pm{.09})$
80.06 (±.$(\pm{.13})$
50.97 (±.$(\pm{.09})$
65.65 (±.$(\pm{.06})$
64.20 (±.$(\pm{.08})$
@@
0.B.2 Cross-Dataset Generalization
@@
From Table 8, we see that our conclusion from Sec 4.3.3 remains valid. Specifically, TPS outperforms TPT [43] by $>2\%$ on average. Similarly to Sec 0.B.1, we observe that taking the mean of the TPT-tuned and advanced off-the-shelf prototypes increases performance by only $0.5\%$ on average over TPT, demonstrating TPT’s inflexibility in utilizing these more robust class representations.
Appendix 0.C Full Ablations
In Sec 0.C.1, we report full ablations on TPS on the effectiveness of feature-space shift on various prototypes. These results are comparable to those reported in Sec 4.5. In Sec 0.C.2, we include additional ablations to observe the effect of learning a class-specific shift over a universal shift for all classes. In Sec 0.C.3, we explore variants on prototype generation using the class-agnostic CLIP ImageNet context prompt templates [38] and the class-specific descriptors generated using GPT-4 [34]. All these ablations are run over 3 random seeds.
0.C.1 Effect of Shift on Different Prototypes
@@
Vanilla
Zero-Shot
66.74
47.79
60.89
73.99
46.12
59.10
57.20
 + shift
68.81 (±.$(\pm{.03})$
58.11 (±.$(\pm{.16})$
63.51 (±.$(\pm{.17})$
76.98 (±.$(\pm{.05})$
48.11 (±.$(\pm{.09})$
63.10 (±.$(\pm{.08})$
61.68 (±.$(\pm{.09})$
$\mathbf{\Delta}$
 + 2.07  + 10.32  + 2.62  + 2.99  + 1.99  + 4.00  + 4.48
CoOp [56]
Zero-Shot
71.51
49.71
64.20
75.21
47.99
61.72
59.28
 + shift
73.76 (±.$(\pm{.04})$
60.43 (±.$(\pm{.12})$
66.84 (±.$(\pm{.10})$
77.39 (±.$(\pm{.05})$
49.08 (±.$(\pm{.06})$
65.50 (±.$(\pm{.02})$
63.44 (±.$(\pm{.03})$
$\mathbf{\Delta}$
 + 2.25  + 10.72  + 2.64  + 2.18  + 1.09  + 3.78  + 4.16
CLIP templates
Zero-Shot
68.35
49.95
61.97
77.59
48.21
61.21
59.43
 + shift
70.39 (±.$(\pm{.06})$
60.47 (±.$(\pm{.07})$
64.66 (±.$(\pm{.04})$
80.70 (±.$(\pm{.04})$
50.38 (±.$(\pm{.14})$
65.32 (±.$(\pm{.03})$
64.05 (±.$(\pm{.02})$
$\mathbf{\Delta}$
 + 2.04  + 10.52  + 2.69  + 3.11  + 2.17  + 4.11  + 4.62
Descriptors
Zero-Shot
68.52
48.91
61.78
74.81
47.68
60.34
58.29
 + shift
70.38 (±.$(\pm{.03})$
59.21 (±.$(\pm{.09})$
63.80 (±.$(\pm{.07})$
77.49 (±.$(\pm{.12})$
49.57 (±.$(\pm{.06})$
64.09 (±.$(\pm{.02})$
62.52 (±.$(\pm{.03})$
$\mathbf{\Delta}$
 + 1.86  + 10.30  + 2.02  + 2.68  + 1.89  + 3.75  + 4.23
CLIP templates
Zero-Shot
69.54
50.51
63.01
77.18
48.84
61.82
59.88
 + Descriptors
 + shift
71.43 (±.$(\pm{.06})$
60.78 (±.$(\pm{.21})$
65.00 (±.$(\pm{.09})$
80.06 (±.$(\pm{.13})$
50.97 (±.$(\pm{.09})$
65.65 (±.$(\pm{.06})$
64.20 (±.$(\pm{.08})$
$\mathbf{\Delta}$
 + 1.89  + 10.27  + 1.99  + 2.88  + 2.13  + 3.83  + 4.32
@@
Full comparisons between zero-shot and feature-shifted performance on all natural distribution shift and cross-domain generalization benchmark datasets over 3 random seeds are in Tables 9 and  10, respectively. We demonstrate that our conclusion from Sec 4.5.1 stills holds – that learning a small perturbation in the feature space results in performance gains of $>4\%$ and up to $1\%$ on average across natural distribution shift and cross-domain generalization tasks regardless of what prototypes are used.
0.C.2 Effect of Per-Class vs. Shared Shift
@@ 0.C.3 Prototype Variants
Vanilla prompt
N/A
66.74
47.79
60.89
73.99
46.12
59.10
57.20
CLIP templates + Descriptors
Macro
68.73
50.32
62.31
77.67
48.56
61.52
59.72
CLIP templates + Descriptors
Micro
69.54
50.51
63.01
77.18
48.84
61.82 59.88
CLIP templates $\times$ Descriptors
Macro
69.03
50.73
62.22
76.91
49.07
61.59
59.73
@@
Vanilla prompt
N/A
68.81 (±.$(\pm{.03})$
58.11 (±.$(\pm{.16})$
63.51 (±.$(\pm{.17})$
76.98 (±.$(\pm{.05})$
48.11 (±.$(\pm{.09})$
63.10 (±.$(\pm{.08})$
61.68 (±.$(\pm{.09})$
CLIP templates + Descriptors
Macro
70.75 (±.$(\pm{.08})$
60.86 (±.$(\pm{.09})$
64.95 (±.$(\pm{.11})$
80.84 (±.$(\pm{.03})$
50.70 (±.$(\pm{.11})$
65.62 (±.$(\pm{.02})$
64.34 (±.$(\pm{.02})$
CLIP templates + Descriptors
Micro
71.43 (±.$(\pm{.06})$
60.78 (±.$(\pm{.21})$
65.00 (±.$(\pm{.09})$
80.06 (±.$(\pm{.13})$
50.97 (±.$(\pm{.09})$
65.65 (±.$(\pm{.06})$
64.20 (±.$(\pm{.08})$
CLIP templates $\times$ Descriptors
Macro
70.82 (±.$(\pm{.02})$
60.42 (±.$(\pm{.06})$
64.50 (±.$(\pm{.05})$
79.53 (±.$(\pm{.09})$
51.13 (±.$(\pm{.02})$
65.28 (±.$(\pm{.01})$
63.89 (±.$(\pm{.02})$
@@
Tables 13 and  14 compare three variants of pooling these CLIP templated embeddings and descriptor embeddings to obtain a single class prototype. Similarly to the conclusion of Sec 4.5.1, we observe that in general, the gains observed using more advanced prototypes in the zero-shot setting almost directly translate to the test-time adaptation setting with shifting. In Sec 4, we present the results of our method using prototypes that are a micro average of the CLIP templates and LLM-generated descriptors.

### v2

@@ 4.1.2 Cross-Dataset Generalization
Ours (Shift + templates + descriptors) 71.54(↑4.26) 50.47(↑6.03) 87.35(↓0.63) 69.06(↑3.82) 71.00(↑5.92) 95.09(↑2.11) 85.23(↑1.43) 68.98(↑6.43) 26.34(↑2.64) 44.48(↑3.06) 66.96(↑3.51)
[…]
@@
4.2 Context-dependent Visual Reasoning
Dataset
We evaluate our method on the Bongard-HOI [23] benchmark, where each test example consists of a small set of 6 positive support images $\mathcal{P}$ illustrating a specific human-object interaction (HOI) concept and 6 negative support images $\mathcal{N}$ that does not demonstrate that concept. The task is to identify whether a given query image $q$ belongs to the positive or negative set.
Implementation Details
We initialize the positive and negative class embeddings from a Gaussian distribution with $\sigma=0.02$, and initialize the learnable shift embedding to all zeros. Similarly to Shu et al. [47], we optimize the shift embedding to minimize the cross-entropy loss on the support set for 64 steps using the AdamW [31] optimizer and learning rate of 5e-3. Then, we classify the query image by cosine similarity of the shifted class embeddings and the query image embedding.
Baselines
We compare our method with fully-supervised and TTA baselines that leverage CLIP-ResNet50 as a backbone. HOITrans [63] is a fully-supervised transformed-based detection method that matches HOIs in the query image to those in the support set. TPT [47] performs text-prompt tuning to encode the HOI concept in the prompt and learn optimal [CLS] embeddings.
4.2.1 Human-Object Interaction
Method seen act., unseen act., seen act., unseen act., Average
seen obj., seen obj., unseen obj., unseen obj.,
Fully-Supervised Baseline
HOITrans [63] 59.50 64.38 63.10 62.87 62.46
Test-Time Adaption Baseline
TPT [47] 66.39 68.50 65.98 65.48 66.59
Ours (Shift) 66.30 72.64 65.63 66.94 67.88
Table 4: Acc on the Bongard-HOI benchmark with CLIP-ResNet-50 backbone. Best performances are in bold.
Table 4 demonstrates the accuracy of our method across the standard test splits of the Bongard-HOI benchmark. We demonstrate improved performance over existing fully-supervised baselines. Notably, we perform better than or on-par with TPT across all test splits, achieving a 1.28% increase in performance, on average.
@@ 4.4.1 Effect of Shift on Different Prototypes
Prompt Type Setting ImageNet ImageNet OOD Cross-Dataset
Average Average
Vanilla Zero-Shot 66.74 57.20 63.45
 + shift 68.77 61.59 64.41
$\mathbf{\Delta}$  + 2.03  + 4.39  + 0.96
CoOp [60] Zero-Shot 71.51 59.28 N/A
 + shift 73.73 63.46 N/A
$\mathbf{\Delta}$  + 2.22  + 4.18 N/A
CLIP templates Zero-Shot 68.35 59.43 64.69
 + shift 70.38 64.04 65.57
$\mathbf{\Delta}$  + 2.03  + 4.61  + 0.88
Descriptors Zero-Shot 68.52 58.29 66.02
 + shift 70.40 62.48 66.80
$\mathbf{\Delta}$  + 1.88  + 4.19  + 0.78
CLIP templates Zero-Shot 69.54 59.88 65.94
+ Descriptors  + shift 71.45 64.15 66.96
$\mathbf{\Delta}$  + 1.91  + 4.27  + 1.02
Table 5: Acc@1 for zero-shot image classification comparing pure zero-shot vs. with learned feature-space shift from prototypes derived from various prompts using CLIP-ViT/B-16 backbone.
@@ 4.4.2 Feature-Space Transformation Variants
Method ImageNet ImageNet OOD Cross-Dataset
Average Average
Scale 70.40 61.12 65.96
@@ 5 Conclusion
We have presented the Test-time Prototype Shifting (TPS) framework, a novel approach to enhancing the zero-shot generalization abilities of VLMs. TPS addresses the limitations of existing test-time training methods by directly modulating class prototypes in the embedding space. This strategy not only reduces the computational and memory demands significantly but also allows for greater flexibility and precision in adapting to diverse domain shifts. By leveraging pre-computed and cached prototypes, and introducing class-specific shifts, TPS effectively bridges domain gaps. The extensive evaluations conducted across many datasets over two visual tasks demonstrate the superior performance of our method, outperforming existing approaches in terms of accuracy while being more computationally efficient.
Acknowledgements This work was supported by the National Science Foundation grant No. 2026498.
Supplementary Material
This document provides more details of our approach and additional experimental results, organized as follows:
  • § A Implementation Details.
  • § B Additional Quantitative Results with Different Random Seeds.
  • § C Additional Ablation Studies.
  • § D Comparison to Training-Free Methods.
  • § E Research Impact and Limitations.
Appendix A Implementation Details of TPS
@@
Appendix B Main Results With More Random Seeds
In Sec B.1 and  B.2, we run Test-Time Prototype Shifting (TPS) over 3 random seeds on both the natural distribution shifts (Table 1) and cross-dataset generalization (Table 3), respectively. The randomness comes from the image augmentation in creating a diverse minibatch for the entropy minimization objective.
B.1 Natural Distribution Shifts
Test-Time Adaptation Baselines
TPT [47] 68.96 (±.$(\pm{.03})$ 54.47 (±.$(\pm{.26})$ 63.46 (±.$(\pm{.07})$ 77.10 (±.$(\pm{.04})$ 47.93 (±.$(\pm{.03})$ 62.38 (±.$(\pm{.05})$ 60.74 (±.$(\pm{.06})$
TPT + (templates + descriptors)* 69.51 (±.$(\pm{.05})$ 54.94 (±.$(\pm{.17})$ 63.86 (±.$(\pm{.11})$ 77.57 (±.$(\pm{.11})$ 48.38 (±.$(\pm{.04})$ 62.85 (±.$(\pm{.03})$ 61.19 (±.$(\pm{.04})$
Ours 71.43 (±.$(\pm{.06})$ 60.78 (±.$(\pm{.21})$ 65.00 (±.$(\pm{.09})$ 80.06 (±.$(\pm{.13})$ 50.97 (±.$(\pm{.09})$ 65.65 (±.$(\pm{.06})$ 64.20 (±.$(\pm{.08})$
@@ B.1 Natural Distribution Shifts
B.2 Cross-Dataset Generalization
@@
From Table 8, we see that our conclusion from Sec 4.1.2 remains valid. Specifically, TPS outperforms TPT [47] by $>2\%$ on average. Similarly to Sec B.1, we observe that taking the mean of the TPT-tuned and advanced off-the-shelf prototypes increases performance by only $0.5\%$ on average over TPT, demonstrating TPT’s inflexibility in utilizing these more robust class representations.
B.3 Context-Dependent Visual Reasoning
Method seen act., unseen act., seen act., unseen act., Average
seen obj., seen obj., unseen obj., unseen obj.,
TPT (reprod.) 65.81 (±.$\pm.12$) 69.15 (±.$\pm.10$) 65.69 (±.$\pm.01$) 66.87 (±.$\pm.03$) 66.88 (±.$\pm.04$)
Ours (Shift) 66.67 (±$\pm 0.68$) 70.31 (±$\pm 1.67$) 66.00 (±$\pm 1.38$) 66.67 (±$\pm 0.40$) 67.41 (±.$\pm.98$)
Table 9: Acc on the Bongard-HOI benchmark with CLIP-ResNet-50 backbone over 3 random seeds. Best performances are in bold.
From Table 9, we see that our conclusion from Sec 4.2 remains valid. Specifically, TPS outperforms TPT [47] by $>0.5\%$ on average.
SuS-X-SD-C [51] 61.65 61.69 35.88 60.49
SuS-X-LC-P [51] 61.80 61.62 36.25 60.64
CALIP [12] 60.57 - - 59.34
Ours (Shift + SuS-X descriptors) 63.52 63.66 37.66 61.47
Appendix C Full Ablations
In Sec C.1, we report full ablations on TPS on the effectiveness of feature-space shift on various prototypes. These results are comparable to those reported in Sec 4.4. In Sec C.2, we include additional ablations to observe the effect of learning a class-specific shift over a universal shift for all classes. In Sec C.3, we explore variants on prototype generation using the class-agnostic CLIP ImageNet context prompt templates [42] and the class-specific descriptors generated using GPT-4 [38]. All these ablations are run over 3 random seeds.
C.1 Effect of Shift on Different Prototypes
Vanilla Zero-Shot 66.74 47.79 60.89 73.99 46.12 59.10 57.20
 + shift 68.81 (±.$(\pm{.03})$ 58.11 (±.$(\pm{.16})$ 63.51 (±.$(\pm{.17})$ 76.98 (±.$(\pm{.05})$ 48.11 (±.$(\pm{.09})$ 63.10 (±.$(\pm{.08})$ 61.68 (±.$(\pm{.09})$
$\mathbf{\Delta}$  + 2.07  + 10.32  + 2.62  + 2.99  + 1.99  + 4.00  + 4.48
CoOp [60] Zero-Shot 71.51 49.71 64.20 75.21 47.99 61.72 59.28
 + shift 73.76 (±.$(\pm{.04})$ 60.43 (±.$(\pm{.12})$ 66.84 (±.$(\pm{.10})$ 77.39 (±.$(\pm{.05})$ 49.08 (±.$(\pm{.06})$ 65.50 (±.$(\pm{.02})$ 63.44 (±.$(\pm{.03})$
$\mathbf{\Delta}$  + 2.25  + 10.72  + 2.64  + 2.18  + 1.09  + 3.78  + 4.16
CLIP templates Zero-Shot 68.35 49.95 61.97 77.59 48.21 61.21 59.43
 + shift 70.39 (±.$(\pm{.06})$ 60.47 (±.$(\pm{.07})$ 64.66 (±.$(\pm{.04})$ 80.70 (±.$(\pm{.04})$ 50.38 (±.$(\pm{.14})$ 65.32 (±.$(\pm{.03})$ 64.05 (±.$(\pm{.02})$
$\mathbf{\Delta}$  + 2.04  + 10.52  + 2.69  + 3.11  + 2.17  + 4.11  + 4.62
Descriptors Zero-Shot 68.52 48.91 61.78 74.81 47.68 60.34 58.29
 + shift 70.38 (±.$(\pm{.03})$ 59.21 (±.$(\pm{.09})$ 63.80 (±.$(\pm{.07})$ 77.49 (±.$(\pm{.12})$ 49.57 (±.$(\pm{.06})$ 64.09 (±.$(\pm{.02})$ 62.52 (±.$(\pm{.03})$
$\mathbf{\Delta}$  + 1.86  + 10.30  + 2.02  + 2.68  + 1.89  + 3.75  + 4.23
CLIP templates Zero-Shot 69.54 50.51 63.01 77.18 48.84 61.82 59.88
+ Descriptors  + shift 71.43 (±.$(\pm{.06})$ 60.78 (±.$(\pm{.21})$ 65.00 (±.$(\pm{.09})$ 80.06 (±.$(\pm{.13})$ 50.97 (±.$(\pm{.09})$ 65.65 (±.$(\pm{.06})$ 64.20 (±.$(\pm{.08})$
$\mathbf{\Delta}$  + 1.89  + 10.27  + 1.99  + 2.88  + 2.13  + 3.83  + 4.32
@@ C.1 Effect of Shift on Different Prototypes
Full comparisons between zero-shot and feature-shifted performance on all natural distribution shift and cross-domain generalization benchmark datasets over 3 random seeds are in Tables 11 and  12, respectively. We demonstrate that our conclusion from Sec 4.4.1 stills holds – that learning a small perturbation in the feature space results in performance gains of $>4\%$ and up to $1\%$ on average across natural distribution shift and cross-domain generalization tasks regardless of what prototypes are used.
C.2 Effect of Per-Class vs. Shared Shift
@@ C.3 Prototype Variants
Vanilla prompt N/A 66.74 47.79 60.89 73.99 46.12 59.10 57.20
CLIP templates + Descriptors Macro 68.73 50.32 62.31 77.67 48.56 61.52 59.72
CLIP templates + Descriptors Micro 69.54 50.51 63.01 77.18 48.84 61.82 59.88
CLIP templates $\times$ Descriptors Macro 69.03 50.73 62.22 76.91 49.07 61.59 59.73
@@
Vanilla prompt N/A 68.81 (±.$(\pm{.03})$ 58.11 (±.$(\pm{.16})$ 63.51 (±.$(\pm{.17})$ 76.98 (±.$(\pm{.05})$ 48.11 (±.$(\pm{.09})$ 63.10 (±.$(\pm{.08})$ 61.68 (±.$(\pm{.09})$
CLIP templates + Descriptors Macro 70.75 (±.$(\pm{.08})$ 60.86 (±.$(\pm{.09})$ 64.95 (±.$(\pm{.11})$ 80.84 (±.$(\pm{.03})$ 50.70 (±.$(\pm{.11})$ 65.62 (±.$(\pm{.02})$ 64.34 (±.$(\pm{.02})$
CLIP templates + Descriptors Micro 71.43 (±.$(\pm{.06})$ 60.78 (±.$(\pm{.21})$ 65.00 (±.$(\pm{.09})$ 80.06 (±.$(\pm{.13})$ 50.97 (±.$(\pm{.09})$ 65.65 (±.$(\pm{.06})$ 64.20 (±.$(\pm{.08})$
CLIP templates $\times$ Descriptors Macro 70.82 (±.$(\pm{.02})$ 60.42 (±.$(\pm{.06})$ 64.50 (±.$(\pm{.05})$ 79.53 (±.$(\pm{.09})$ 51.13 (±.$(\pm{.02})$ 65.28 (±.$(\pm{.01})$ 63.89 (±.$(\pm{.02})$
@@
Tables 15 and  16 compare three variants of pooling these CLIP templated embeddings and descriptor embeddings to obtain a single class prototype. Similarly to the conclusion of Sec 4.4.1, we observe that in general, the gains observed using more advanced prototypes in the zero-shot setting almost directly translate to the test-time adaptation setting with shifting. In Sec 4, we present the results of our method using prototypes that are a micro average of the CLIP templates and LLM-generated descriptors.
Appendix D Comparison to Training-Free Methods
Appendix E Research Impact and Limitations